OpenStudy (anonymous):

find the number that makes the ratios equvilent: 8/9=___/36 15/18=5:___ ___ to 7 = 9 to 21

OpenStudy (mokeira):

\[\frac{ 8 }{9 }= \frac{ x }{ 36 }\] cross multiply 8x36=9x 288=9x divide both sides by 9 to get the value you are looking for

OpenStudy (texaschic101):

ratios are equivalent fractions 8/9 and x/36.....what equation is equivalent to 8/9 and has a denominator of 36. To find an equivalent fraction, you have to multiply the given fraction by the same number...for example : 2/4 = x/8....4 times what is 8 ? 4 x 2 = 8. And if the denominator is multiplied by 2, then the numerator has to be multiplied by 2 as well. giving us : 2/4 * 2/2 = 4/8......so 2/4 and 4/8 are equivalent fractions. You see how I multiplied the numerator and denominator by the same number. so 8/9 = x/36.....9 times what is 36 ? 9 x 4 = 36. So if we had to multiply the denominator by 4, then we have to multiply the numerator by 4. 8/9 * 4/4 = 32/36.....so the equivalent fraction of 8/9 with a denominator of 36 is 32/36 . As far as this one...15/18 = 5: x.......keep in mind that ratios can also be written as fractions and vice-versa......for example 15/18 is also 15 : 18. If you can reduce the fraction, do so. 15/18 reduces to 5/6, which is the same as 5:6. and to see if they are equivalent, see if the numerator and denominator are multiplied by the same number. 5/6 * 3/3 = 15/18.....so 15/18 is equivalent to 5/6 or 5:6 As far as the lat one.....I will let you do that one....but keep in mind ratios can be written as fractions. for example : 2 to 4 = 2/4 = 2:4.....all the same